OpenStudy (anonymous):

Is the function quadratic? Explain. y + 5 = 2x^2

OpenStudy (anonymous):

Check for the highest power in this equation.. A Quadratic Function is that in which highest power also called Degree is 2..

jimthompson5910 (jim_thompson5910):

Do you think it's a quadratic? Yes? No?

OpenStudy (anonymous):

Huh ? @waterineyes & Yes ?? I'm not sure.

jimthompson5910 (jim_thompson5910):

well we have an x^2 term so that's a big indication that it is since there are no x^3 terms, x^4 terms, etc this confirms that it is indeed a quadratic

jimthompson5910 (jim_thompson5910):

the rule is: if there is an x^2 term and there are no other terms with larger exponents, then it's a quadratic

jimthompson5910 (jim_thompson5910):

so something like y = x^2 + 5 is a quadratic but something like y = x^2 + 5x^3 is not

OpenStudy (anonymous):

So if it a Quadratic that mean it would beable to be written in the form y = ax2 + bx + c form ?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

okay(: thank you

jimthompson5910 (jim_thompson5910):

y + 5 = 2x^2 y = 2x^2 - 5 y = 2x^2 + 0x + (-5) which is now in the form y = ax^2 + bx + c where a=2, b = 0 and c = -5

jimthompson5910 (jim_thompson5910):

yw